OpenStudy (lexber):

will give medal and fan(: Write the equation of the line that passes through the points (3, 6) and (5, 18) using function notation. f(x) = 2x + 12 f(x) = 6x − 12 y = 2x + 12 y = 6x − 12

OpenStudy (isaiahjwolfe):

This question wants you to fill in the function f(x) = mx - b

OpenStudy (isaiahjwolfe):

m stands for the slope. To find the slope you use the average rate of change formula:\[\frac{ y2-y1 }{ x2 - x1 }\]

OpenStudy (isaiahjwolfe):

You fill it in with the points given to you in the question: (3,6) and (5,18). Can you fill the formula in? (If not just tell me and I will walk you through it)

OpenStudy (isaiahjwolfe):

@lexber

OpenStudy (lexber):

no i cant i need help @isaiahjwolfe

OpenStudy (isaiahjwolfe):

Okay, let's take (3,6). The x and y in a point is (x,y). So in (3,6) x = 3, and 6 = y. Let's insert these into our formula:\[\frac{ y2 - 6 }{ x2 - 3 }\]

OpenStudy (lexber):

okay

OpenStudy (isaiahjwolfe):

Know that I showed you how to insert one point, can you insert (5,18) into the formula?

OpenStudy (isaiahjwolfe):

Now*

OpenStudy (lexber):

umm

OpenStudy (lexber):

idk this is hard for me im really bad at math

OpenStudy (isaiahjwolfe):

Can you identify the x and y in (5,18)?

OpenStudy (lexber):

yes i think

OpenStudy (isaiahjwolfe):

What are they

OpenStudy (lexber):

y2=5 and x2=18

OpenStudy (isaiahjwolfe):

Yes!

OpenStudy (lexber):

yay!

OpenStudy (isaiahjwolfe):

So now let's insert them into the formula:\[\frac{ 18 - 6 }{ 5 - 3 }\]Can you solve it?

OpenStudy (lexber):

6/2 ?

OpenStudy (isaiahjwolfe):

Yup, I think B will be your answer

OpenStudy (isaiahjwolfe):

I have to go, I hope you do well on your quiz :)